# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ III-C5 R-NaD
Although R-NaD [54] is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a special oracle computation technique (Algo. 3), which is executed in two stages: the first stage involves transforming the reward based on the regulation policy to make it policy-dependent, and the second stage applies replicator dynamics [32] until convergence to a fixed point. Crucially, in each iteration, the oracle added to the policy population $\Pi$ is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy will converge to the NE of the original game, provided that the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (33), the same as the MSS of vanilla self-play. This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
III-D Ongoing-training-based Series of Algorithms
@@
In the PSRO series of algorithms, two key challenges arise. First, when operating with a limited budget, it is often necessary to truncate the ABR operators during each iteration. This can introduce sub-optimally trained responses into the population. Second, the redundant process of relearning basic skills in every iteration is not only inefficient, but it also becomes untenable when confronted with increasingly formidable opponents [43]. To address these challenges, the ongoing-training-based series of algorithms promote the ongoing training of all policies repeatedly. Namely, all effective policies are likely to be selected for training.
III-D1 Integration into Our Framework
@@
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use actual initialization to initialize $\Pi$ because in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for a total of $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$. This means that policy updates are self-referential.
For the sake of simplicity, we also adopt Assumption 1. Different from Collory 1 and Collory 2, due to the continuous training process of all policies, we derive Collory 3.
@@
When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an effective policy. Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning. As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
III-D2 FTW
Quake III Arena Capture the Flag is a renowned 3D multiplayer first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent [70] is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
$\mathcal{F}(\mathcal{P})_{ij}=\frac{1}{N}.$ (38)
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$. Furthermore, since these policies are not conditioned on any external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
III-D3 NeuPL
NeuPL [43] introduces another key innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$. Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
III-D4 Simplex-NeuPL
@@
III-E Regret-minimization-based Series of Algorithms
Another line of self-play algorithms is based on regret minimization. The key distinction between regret-minimization-based algorithms and the other categories is that they prioritize accumulated payoff over time, rather than focusing solely on a single episode. This approach leads to more aggressive and adaptable strategies, which are essential to avoid being exploited by opponents as time progresses. Additionally, these algorithms require players to deduce and adapt to opponents’ strategies over several rounds. This scenario is commonly observed in repeated games rather than stage games. For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game. It’s important to note that while traditional regret-minimization-based self-play doesn’t typically use RL, many subsequent research efforts have combined regret minimization with RL to achieve strong performance. In this section, we will also thoroughly discuss traditional regret-minimization-based methods to provide a foundation for understanding how integrating regret minimization with RL can lead to enhanced performance.
III-E1 Integration into Our Framework
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use placeholder initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy. Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy. More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$. Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. It’s important to note that the regret-minimization-based series relies heavily on the information contained in $h(i)$. For instance, in vanilla Counterfactual Regret Minimization (CFR) [72], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$ and once $h(i)$ is determined, the corresponding policy is also defined through regret matching. We will discuss vanilla CFR in detail in Sec. III-E2. Fifth, the ABR operator is described in Algo. 4, with the key point being to incorporate new regret minimization-based information into $h(i)$. Notably, while the original CFR updates regrets for all players simultaneously, this oracle (Algo. 4) updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. Namely, $h(i)$ is changing during the iteration $i$. This adjustment has not only been shown to empirically accelerate convergence but also possesses a theoretical error bound [73]. Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
Moreover, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (33). We can derive Collory 4.
@@ III-E1 Integration into Our Framework
Regret minimization-based algorithms always use the latest strategy for training. In other words, in iteration $i$, $\pi_{i-1}^{h}$ is always chosen as the opponent policy. Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix. ∎
III-E2 Vanilla CFR
Regret measures the difference between the actual payoff and the best possible payoff that could have been achieved with a different strategy. The regret matching algorithm [74] in game theory optimizes decisions in iterative games by learning from past outcomes. More specifically, it involves selecting strategies based on accumulated positive overall regrets. Strategies with higher overall regret are generally more likely to be chosen, as players seek to correct past underperformance. After each round, players update the overall regret values for each strategy. When each player aims to reduce their average overall regret, their average strategies will converge towards an approximation of the NE over time. However, this traditional regret minimization algorithm applies primarily to normal-form games because computing overall regret in extensive-form games is challenging. Although extensive-form games can theoretically be converted into normal-form games, this conversion results in an exponential increase in states, rendering it impractical for complex extensive-form games.
[72] proposes counterfactual regret (CFR) minimization for extensive-form games with incomplete information. Here, it is referred to as vanilla CFR to distinguish it from subsequent advancements in the field. Vanilla CFR involves maintaining the strategy and the counterfactual regrets for each information set. Theoretically, the immediate counterfactual regret is specific to each information set, and the aggregation of these immediate counterfactual regrets can act as an upper bound for the overall regret. Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set. This simplification significantly reduces the computational complexity.
Next, there is a more detailed description of vanilla CFR. In this survey, $\pi_{i}$ is used to denote the strategy at iteration $i$, whereas the original paper uses $\sigma$. This change is implemented to maintain consistency throughout the discussion in this survey. Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$. The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
@@ III-E2 Vanilla CFR
Therefore, based on the theory discussed, the overall problem of regret minimization can be decomposed into numerous smaller regret minimization problems. This approach makes the problem manageable for extensive-form games of not excessively large sizes. In practical applications, the focus is primarily on immediate counterfactual regrets. For simplicity, we often drop "immediate" in discussions, thereby referring directly to counterfactual regrets. Consequently, the counterfactual regret associated with every action $a$ within every information set $I$ is recorded as follows:
@@
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, which becomes computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$. These values are stored in $h(i)$ within our proposed framework (Algo. 1). This requirement leads to significant storage challenges.
III-E3 Time-saving Variants of CFR
Therefore, numerous studies focus on enhancing the time efficiency of CFR primarily through two main approaches. The first approach involves modifying the regret calculation to increase its speed. For example, CFR+ [75] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$. Note that $R_{j}^{+,i}(I,a)$ is computed slightly differently from $R_{j}^{i,+}(I,a)$ in Equ. (45). Additionally, CFR+ updates the regrets of each player sequentially. Furthermore, CFR+ adopts a weighted average strategy instead of a uniform average strategy, assigning a lower weight to the initial strategy for accelerated convergence. Beyond employing a weighted average strategy, [76] introduces the concept of weighted regrets and developed Discounted CFR (DCFR). Linear CFR (LCFR) is a specific case of DCFR, where LCFR assigns a linear weight of $i$ to regrets at iteration $i$.
@@ III-E3 Time-saving Variants of CFR
The other approach involves adopting sampling methods. Although this approach requires more iterations, the duration of each iteration is reduced, ultimately decreasing the overall time to convergence. Monte Carlo CFR (MCCFR) [77] outlines a framework to incorporate sampling into CFR. This method divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for the calculation of sampled counterfactual values for each player, leading to immediate counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR represents a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, focusing on chance nodes alone. Moreover, [78] expands on chance-sampling into naive chance sampling, opponent/public chance sampling, self/public chance sampling, and public chance sampling. These methods differ in how they handle the sampling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [79, 80].
In addition to these two primary approaches, other studies have identified that warm starting [81] and pruning techniques [82, 83, 84] can accelerate CFR convergence as well.
III-E4 Space-saving Variants of CFR
@@
Storing the strategies and regrets of each information set at each iteration requires substantial storage capacity. In games with perfect information, decomposition is employed to reduce the scale of problem-solving; that is, we solve subgames instead of the entire game. However, this approach faces challenges in imperfect information games. The difficulty arises in defining subgames, as they may intersect with the boundaries of information sets, complicating their delineation. CFR-D [85] is a pioneering method for decomposing imperfect information games with a theoretical guarantee. The game is divided into a main component, called the trunk, and several subgames. It posits that in an imperfect information game, a subgame can be conceptualized as a forest of trees that does not divide any information sets. In each iteration, CFR is applied within the trunk for both players, accompanied by the use of a solver to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root. Following these updates, the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving values $R^{i}(I_{*},a)$ for information sets located in the trunk and at each subgame’s root, which is denoted by $I_{*}$ at each iteration $i$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [12] and the Safe and Nested Subgame Solving technique used by Libratus [86]. We will discuss these approaches in Sec. IV-B1, where we explore their application to the specific context of Texas Hold’em.
III-E5 Estimation Variants of CFR
Although the above variants of CFR advance the field, they still fail to solve large imperfect-information extensive-form games directly. This limitation largely stems from the reliance on tabular representations for strategies. Typically, the process involves abstracting the original large game into a simpler form, applying the above CFR variants to this abstracted game, and then translating the developed strategies back to the original game. However, this abstraction process is highly game-specific and is heavily based on domain knowledge. Moreover, abstracting the game to a smaller scale often leads to suboptimal results compared to a more comprehensive abstraction, making the choice of abstraction scale critical for performance. Given these challenges, there is a shift towards estimation techniques. [87] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, the use of regression trees as the regressor limits RCFR’s applicability to smaller games, and the necessity for manually crafted features remains a drawback.

### v2

@@ 3.3.5 R-NaD
Although R-NaD [105] is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo. 3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics [124] until convergence to a fixed point. The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (5). This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
3.4 Ongoing-training-based Series
@@
In the PSRO series of algorithms, two key challenges arise. First, truncating the ABR operators during each iteration is often necessary when operating with a limited budget, introducing sub-optimally trained responses into the population. Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents [82]. To address these challenges, the ongoing-training-based series of algorithms emphasizes the repeated ongoing training of all policies. In other words, instead of only inactive policies being selected for training, all active policies are also likely to be chosen for training.
3.4.1 Integration into Our Framework
@@
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use immediate initialization to initialize $\Pi$ because, in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$. This means that policy updates are self-referential. Lastly, under Assumption 1, different from Collory 1 and 2, due to the continuous training process of all policies, we derive Collory 3.
@@
When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an active policy. Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning. As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
3.4.2 FTW
Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent [60] is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
$\textit{MSS}(\mathcal{P})_{ij}=\frac{1}{N}.$ (10)
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$. Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
3.4.3 NeuPL
Neural Population Learning (NeuPL) [82] introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$. Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
3.4.4 Simplex-NeuPL
@@
3.5 Regret-minimization-based Series
Another line of self-play algorithms is based on regret minimization. Unlike other categories, this series prioritizes accumulated payoff over time rather than focusing solely on a single episode, which is common in repeated games. For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game. Traditional regret-minimization-based self-play typically doesn’t use RL, but later studies have combined them for improved performance. This section will also discuss traditional regret-minimization methods to lay the groundwork for understanding their integration with RL.
3.5.1 Integration into Our Framework
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use lazy initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy. Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy. More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$. Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. Note that this category relies heavily on the information in $h(i)$. For instance, in Counterfactual Regret Minimization (CFR) [152], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$. Once $h(i)$ is determined, the corresponding policy is also defined. We will discuss CFR in detail in Sec. 3.5.2. Fifth, the ABR operator (Algo. 4) incorporates regret-related information into $h(i)$. Unlike the original CFR, Algo. 4 updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound [22]. Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$. Lastly, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (5). Also, we can derive Collory 4.
@@ 3.5.1 Integration into Our Framework
Regret minimization-based algorithms always use the latest strategy for training. In other words, at iteration $i$, $\pi_{i-1}^{h}$ is consistently chosen as the opponent policy. Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix. ∎
3.5.2 Vanilla CFR
Regret measures the difference between the best possible payoff and the actual payoff. The regret matching algorithm [49] optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance. After each round, players update the overall regret values for each strategy. Strategies will converge to an approximate NE. However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
[152] proposes CFR for extensive-form games. Here, it is referred to as vanilla CFR to distinguish it from later advancements. Vanilla CFR maintains the strategy and the immediate counterfactual regrets for each information set. Theoretically, this regret is specific to each information set, and the aggregation of these immediate counterfactual regrets is an upper bound for the overall regret. Therefore, the problem can be decomposed into numerous more minor regret minimization problems.
Next, there is a more detailed description of vanilla CFR. In this survey, $\pi_{i}$ denotes the strategy at iteration $i$ for consistency, whereas the original paper uses $\sigma$. Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$. The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
@@ 3.5.2 Vanilla CFR
Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set separately. For simplicity, we often drop "immediate" in discussions and refer directly to counterfactual regrets. The counterfactual regret can also be recorded as follows:
@@
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$. These values are stored in $h(i)$ within our proposed framework (Algo. 1), leading to significant storage challenges.
3.5.3 Time-saving Variants of CFR
@@
Many studies focus on enhancing the time efficiency of CFR. The first approach involves modifying the regret calculation to increase its speed. CFR+ [122] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$. Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy. Moreover, [19] introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR). The second approach involves adopting sampling methods. While it requires more iterations, each iteration is shorter, reducing the overall convergence time. Monte Carlo CFR (MCCFR) [72] divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for calculating sampled counterfactual values for each player, leading to counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR is a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, only focusing on chance nodes. Moreover, [64] expands on chance-sampling by categorizing based on the handling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [111, 33]. In addition to these two primary approaches, other studies have identified that warm starting [16] and pruning [15, 13, 17] can also accelerate convergence.
3.5.4 Space-saving Variants of CFR
@@
In perfect information games, decomposition reduces problem-solving scale by solving subgames. However, in imperfect information games, defining subgames is challenging due to their intersection with information set boundaries. CFR-D [21] is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets. In each iteration, CFR is applied within the trunk for both players, and a solver is used to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root, while the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving $R^{i}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [95] and the Safe and Nested Subgame Solving technique used by Libratus [18]. We will discuss these approaches in Sec. 4.2.1, exploring their application to Texas Hold’em.
3.5.5 Estimation Variants of CFR